# Partial Law Invariance and Risk Measures

### v4

@@ 3.3 Two constructions
There is a large literature on conditional and dynamic risk measures, which also involve a fixed sub-$\sigma$-algebra $\mathcal{G}$; see Riedel (2004), Detlefsen and Scandolo (2005), and Filipović et al. (2012). A conditional risk measures is a mapping $\rho:L^{\infty}\to L^{\infty}(\mathcal{G})$ satisfying some specific properties. We emphasize that this paper deals with the static setting because our risk measures take real values. However, it is possible, as shown above, to use conditional coherent risk measures to construct $\mathcal{G}$-law-invariant coherent risk measures, but our mathematical results do not overlap with the results in the above literature. Moreover, the motivation for partial law invariance differs from that of dynamic risk measures, although both involve sub-$\sigma$-algebras.
@@
Functionals of the form (6) can be found in the two-stage market evaluation of Pelsser and Stadje (2014) and the market-consistent valuation of Dhaene et al. (2017). Setting $\tau(X)=\mathbb{E}[X|\mathcal{G}]$, equation (6) offers a straightforward method for constructing partially law-invariant risk measures. Specifically,
@@
The risk measure in (7) further satisfies a stronger form of partial law invariance, which will be studied in Section 5.2.
Example 3.
The idea of using conditional expectations can be generalized further to incorporate different sub-$\sigma$-fields (see Section 4). Let $\mathcal{H}$ be a sub-$\sigma$-algebra independent of $\mathcal{G}$. Let $\rho_{1}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ and $\rho_{2}:L^{\infty}(\mathcal{H})\to\mathbb{R}$ both be coherent risk measures, with the former being law invariant. Define the following risk measure
@@
The second class of partially law-invariant coherent risk measures arises from statistical functions and tail risk measures, similar to the example of robust ES in Section 2.2.
Let $\mathcal{P}_{c}(\mathbb{R})$ denote the set of Borel probability measures on $\mathbb{R}$ with compact support. A statistical function is mapping $\gamma:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$. We say that a statistical function is coherent if the functional on $L^{\infty}$ given by $X\mapsto\gamma(X_{\#}\mathbb{P})$ is coherent, where $X_{\#}\mathbb{P}\in\mathcal{P}_{c}(\mathbb{R})$ is the distribution of $X$ under $\mathbb{P}$. We say that a statistical function is Fatou if the functional on $L^{\infty}$ given by $X\mapsto\gamma(X_{\#}\mathbb{P})$ is Fatou.
Proposition 7.
@@
$\rho(X)=\sup_{\mu\in\mathcal{S}}\gamma(X_{\#}\mu),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}$ (8)
@@
Using Proposition 7, we can establish a connection between a “bottom-up” approach to addressing model uncertainty and the current framework. Suppose that an agent begins with a probability space $(\Omega,\mathcal{G},\mathbb{P}_{0})$, where they are certain about their probabilistic model $\mathbb{P}_{0}$ on $\mathcal{G}$. The agent wishes to extend this model to a finer-$\sigma$-algebra $\mathcal{F}\supseteq\mathcal{G}$, on which they have model uncertainty. To start, she defines a reference probability $\mathbb{P}$ on $\mathcal{F}$, such that $\mathbb{P}|_{\mathcal{G}}=\mathbb{P}_{0}$, to identify null sets. From here, the agent chooses a set of competing probabilistic models $\mathcal{S}\subseteq\mathcal{E}(\mathbb{P}_{0})$. For financial positions in $L^{\infty}(\mathcal{G})$, the agent intended to use the law-invariant risk measure $\rho:L^{\infty}(\mathcal{G})\to\mathbb{R}$ given by $\tilde{\rho}(X)=\gamma(X_{\#}\mathbb{P}_{0}),$ $X\in L^{\infty}(\mathcal{G})$, where $\gamma$ is a statistical function that is Fatou and coherent. The risk measure defined in equation (8) would be a natural risk measure to use on $L^{\infty}$. According to Proposition 7, this risk measure is Fatou, $\mathcal{G}$-law invariant, and coherent. For a similar approach in the dynamical setting, see Epstein and Schneider (2003) and Hansen and Sargent (2022).
Example 4.
For $\beta\in[0,1)$, define $\mathcal{S}=\{\mu\in\mathcal{E}(\mathbb{P}):\leavevmode\nobreak\ D_{\mu}\leq 1/(1-\beta)\}$ and let $\tilde{\mathbb{E}}:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ be the expectation statistical function (mapping a distribution to its mean). Define $\rho:L^{\infty}\to\mathbb{R}$ by
$\rho(X)=\sup_{\mu\in\mathcal{S}}\tilde{\mathbb{E}}(X_{\#}\mu)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
In this case, $\rho$ is of the form in (4) and (8). As $\mathcal{S}^{\mathcal{G}}=\{\mathbb{P}|_{\mathcal{G}}\}$ is almost $\mathcal{G}$-law invariant, we can also verify that $\rho$ is $\mathcal{G}$-law invariant from Theorem 1. Take any $\sigma$-algebra $\mathcal{H}$ independent of $\mathcal{G}$, such that $(\Omega,\mathcal{H},\mathbb{P})$ is atomless (assumed to exist; this corresponds to a notion of conditional atomless property; see Remark 4 in Appendix B). We claim
$\rho(X)=\mathbb{E}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{G})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \rho(X)=\mathrm{ES}_{\beta}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ Y\in L^{\infty}(\mathcal{H}).$ (9)
@@
where $\tilde{\mathrm{ES}}_{\beta}:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ is the ES statistical function. The $\sigma$-algebras $\mathcal{G}$ and $\mathcal{H}$ are extreme as $\rho(X)=\tilde{\mathbb{E}}(Q)$ if $X\in L^{\infty}(\mathcal{G})$ and $\rho(X)=\tilde{\mathrm{ES}}_{\beta}(Q)$ if $X\in L^{\infty}(\mathcal{H})$. For proofs of the above claims, see Section C.3.
@@
By allowing $\mathcal{S}$ to vary but maintaining the condition that $\mathcal{S}^{\mathcal{G}}=\{\mathbb{P}|_{\mathcal{G}}\}$, we allow the risk measure to behave “nicely” on $L^{\infty}(\mathcal{G})$, while leaving flexibility outside this $\sigma$-field. Certainly, it is not necessary to require $\mu\in\mathcal{E}(\mathbb{P})$ for all $\mu\in\mathcal{S}$ (like $\mathcal{S}$ in Example 4); according to Theorem 1, it is necessary and sufficient to take $\mathcal{S}^{\mathcal{G}}$ to be almost $\mathcal{G}$-law invariant. Indeed, generalizing Example 4, we can have $\rho$ behave like a law-invariant risk measure $\widehat{\rho}$ on $L^{\infty}(\mathcal{G})$ and like a tail risk measure generated by $\widehat{\rho}$ on $L^{\infty}(\mathcal{H})$ in the sense of Liu and Wang (2021), this is illustrated by Example 5.
Example 5.
@@
Given a Fatou law-invariant coherent risk measure $\widehat{\rho}:L^{\infty}\to\mathbb{R}$, there exists a Fatou coherent statistical function $\gamma:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ such that $\widehat{\rho}(X)=\gamma(X_{\#}\mathbb{P})$ for all $X\in L^{\infty}$. For $\beta\in[0,1)$, define $\mathcal{S}=\{\mu\in\mathcal{E}(\mathbb{P}):\leavevmode\nobreak\ D_{\mu}\leq 1/(1-\beta)\}$. By Proposition 7, the risk measure $\rho$ given by (8) is Fatou, $\mathcal{G}$-law invariant, and coherent. We claim that the risk measure $\rho$ satisfies
$\rho(X)=\widehat{\rho}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{G})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \rho(X)=\widehat{\rho}^{\beta}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{H}),$

### v5

@@ 6.2 Two constructions of partially law-invariant coherent risk measures
There is a large literature on conditional and dynamic risk measures, which also involve a fixed sub-$\sigma$-algebra $\mathcal{G}$; see Riedel (2004), Detlefsen and Scandolo (2005), and Filipović et al. (2012). Conditional risk measures take values in random variables, different from our setting, and the motivation for partial law invariance is quite different from the above literature. Nevertheless, it is possible, as shown above, to use the theory of conditional coherent risk measures to construct $\mathcal{G}$-law-invariant coherent risk measures.
@@
Functionals of the form (12) can be found in the two-stage market evaluation of Pelsser and Stadje (2014) and the market-consistent valuation of Dhaene et al. (2017). Setting $\psi(X)=\mathbb{E}[X|\mathcal{G}]$, equation (12) offers a straightforward method for constructing partially law-invariant risk measures. Specifically,
@@
Example 6.6
The idea of using conditional expectations can be generalized further to incorporate different sub-$\sigma$-algebras (see Section 4). Let $\mathcal{H}$ be a sub-$\sigma$-algebra independent of $\mathcal{G}$. Let $\rho_{1}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ and $\rho_{2}:L^{\infty}(\mathcal{H})\to\mathbb{R}$ both be coherent risk measures, with the former being law invariant. Define the following risk measure
@@
The second class of partially law-invariant coherent risk measures arises from statistical functions and tail risk measures, similar to the example of robust ES.
Let $\mathcal{P}_{c}(\mathbb{R})$ denote the set of Borel probability measures on $\mathbb{R}$ with compact support. A statistical function is a mapping $\gamma:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$. We say that a statistical function is coherent if the functional on $L^{\infty}$ given by $X\mapsto\gamma(X_{\#}\mathbb{P})$ is coherent, where $X_{\#}\mathbb{P}\in\mathcal{P}_{c}(\mathbb{R})$ is the distribution of $X$ under $\mathbb{P}$. We say that a statistical function is Fatou if the functional on $L^{\infty}$ given by $X\mapsto\gamma(X_{\#}\mathbb{P})$ is Fatou.
Proposition 6.7
@@
$\rho(X)=\sup_{\mu\in\mathcal{S}}\gamma(X_{\#}\mu),~{}~{}~{}X\in L^{\infty}$ (14)
@@
Using Proposition 6.7, we can establish a connection between a “bottom-up” approach to addressing model uncertainty and the current framework. Suppose that an agent begins with a probability space $(\Omega,\mathcal{G},\mathbb{P}_{0})$, where they are certain about their probabilistic model $\mathbb{P}_{0}$ on $\mathcal{G}$. The agent wishes to extend this model to a finer-$\sigma$-algebra $\mathcal{F}\supseteq\mathcal{G}$, on which they have model uncertainty. To start, she defines a reference probability $\mathbb{P}$ on $\mathcal{F}$, such that $\mathbb{P}|_{\mathcal{G}}=\mathbb{P}_{0}$, to identify null sets. From here, the agent chooses a set of competing probabilistic models $\mathcal{S}\subseteq\mathcal{E}(\mathbb{P}_{0})$. For financial positions in $L^{\infty}(\mathcal{G})$, the agent intends to use the law-invariant risk measure $\rho:L^{\infty}(\mathcal{G})\to\mathbb{R}$ given by $\tilde{\rho}(X)=\gamma(X_{\#}\mathbb{P}_{0}),$ $X\in L^{\infty}(\mathcal{G})$, where $\gamma$ is a statistical function that is Fatou and coherent. The risk measure defined in equation (14) would be a natural risk measure to use on $L^{\infty}$. According to Proposition 6.7, this risk measure is Fatou, $\mathcal{G}$-law invariant, and coherent. For a similar approach in the dynamical setting, see Epstein and Schneider (2003) and Hansen and Sargent (2022).
Example 6.8
For $\beta\in[0,1)$, define $\mathcal{S}=\{\mu\in\mathcal{E}(\mathbb{P}):~{}D_{\mu}\leq 1/(1-\beta)\}$ and let $\tilde{\mathbb{E}}:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ be the expectation statistical function (mapping a distribution to its mean). Define $\rho:L^{\infty}\to\mathbb{R}$ by
$\rho(X)=\sup_{\mu\in\mathcal{S}}\tilde{\mathbb{E}}(X_{\#}\mu)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],~{}~{}X\in L^{\infty}.$
In this case, $\rho$ is of the form in (4) and (14). As $\mathcal{S}^{\mathcal{G}}=\{\mathbb{P}|_{\mathcal{G}}\}$ is almost $\mathcal{G}$-law invariant, we can also verify that $\rho$ is $\mathcal{G}$-law invariant from Theorem 3.3. Take any $\sigma$-algebra $\mathcal{H}$ independent of $\mathcal{G}$, such that $(\Omega,\mathcal{H},\mathbb{P})$ is atomless (assumed to exist; this corresponds to a notion of conditional atomless property; see Remark 9.4 in Appendix 9). We claim
$\rho(X)=\mathbb{E}[X],~{}~{}X\in L^{\infty}(\mathcal{G})~{}~{}\text{and}~{}~{}\rho(X)=\mathrm{ES}_{\beta}(X),~{}~{}Y\in L^{\infty}(\mathcal{H}).$ (15)
@@
where $\tilde{\mathrm{ES}}_{\beta}:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ is the ES statistical function. The $\sigma$-algebras $\mathcal{G}$ and $\mathcal{H}$ are extreme as $\rho(X)=\tilde{\mathbb{E}}(Q)$ if $X\in L^{\infty}(\mathcal{G})$ and $\rho(X)=\tilde{\mathrm{ES}}_{\beta}(Q)$ if $X\in L^{\infty}(\mathcal{H})$. For proofs of the above claims, see Section 13.2.
@@
By allowing $\mathcal{S}$ to vary but maintaining the condition that $\mathcal{S}^{\mathcal{G}}=\{\mathbb{P}|_{\mathcal{G}}\}$, we allow the risk measure to behave “nicely” on $L^{\infty}(\mathcal{G})$, while leaving flexibility outside this $\sigma$-field. Certainly, it is not necessary to require $\mu\in\mathcal{E}(\mathbb{P})$ for all $\mu\in\mathcal{S}$ (like $\mathcal{S}$ in Example 6.8); according to Theorem 3.3, it is necessary and sufficient to take $\mathcal{S}^{\mathcal{G}}$ to be almost $\mathcal{G}$-law invariant. Indeed, generalizing Example 6.8, we can have $\rho$ behave like a law-invariant risk measure $\widehat{\rho}$ on $L^{\infty}(\mathcal{G})$ and like a tail risk measure generated by $\widehat{\rho}$ on $L^{\infty}(\mathcal{H})$ in the sense of Liu and Wang (2021), this is illustrated by Example 6.9.
Example 6.9
@@
Given a Fatou law-invariant coherent risk measure $\widehat{\rho}:L^{\infty}\to\mathbb{R}$, there exists a Fatou coherent statistical function $\gamma:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ such that $\widehat{\rho}(X)=\gamma(X_{\#}\mathbb{P})$ for all $X\in L^{\infty}$. For $\beta\in[0,1)$, define $\mathcal{S}=\{\mu\in\mathcal{E}(\mathbb{P}):~{}D_{\mu}\leq 1/(1-\beta)\}$. By Proposition 6.7, the risk measure $\rho$ given by (14) is Fatou, $\mathcal{G}$-law invariant, and coherent. We claim that the risk measure $\rho$ satisfies
$\rho(X)=\widehat{\rho}(X),~{}~{}X\in L^{\infty}(\mathcal{G})~{}~{}~{}\text{and}~{}~{}~{}\rho(X)=\widehat{\rho}^{\beta}(X),~{}~{}X\in L^{\infty}(\mathcal{H}),$